# A subset $S$ of vertices in a planar graph $G$ is a free set if, for every set $P$ of $|S|$ points in the plane, there exists a straight-line crossing-free drawing of $G$ in which vertices of $S$ are mapped to distinct points in $P$. In this survey, we review

### v1

@@ 5 A One-Bend Variant
A $k$-bend crossing-free drawing of a planar graph $G$ is a crossing-free drawing of $G$ in which each edge is represented by a polygonal chain consisting of at most $k$ line segments. Thus, a 00-bend crossing-free drawing is a straight-line crossing-free drawing. This leads naturally to a definition of a $k$-bend free set introduced by Bose et al. [16]. An ordered set $S:=(v_{1},\ldots,v_{s})$ of vertices in a planar graph $G$ is a $k$-bend free set if, for any $x_{1}<\cdots<x_{s}$ and any $y_{1},\ldots,y_{s}$, there exists a $k$-bend crossing-free drawing $\Gamma$ of $G$ such that $(x_{i},y_{i})$ is the location of $v_{i}$ in $\Gamma$, for each $i\in\{1,\ldots,s\}$. As an application of 9, below, Bose et al. [16] prove the following result:
@@
They prove 8 by proving the following result, which the main result in [16]:
@@
Refer to Fig. 12. Fix some straight-line crossing-free drawing of $G$, which also fixes a straight-line crossing-free drawing of $T$. Fatten the drawing of $T$ by some arbitrarily small value $\epsilon>0$ by taking the Minkowsky sum of $T$ with a disc of radius $\epsilon$. The boundary of this fattened tree is a simple closed curve $C_{0}$. Construct a curve $C$ by deforming $C_{0}$ in an $\epsilon$-neighbourhood of each leaf $v$ of $T$ so that it contains $v$. If $\epsilon$ is sufficiently small, then $C$ is a simple closed curve that intersects each edge $vw$ of $G$ in at most $2$ points: one point within distance $\epsilon$ of $v$ and one point within distance $\epsilon$ of $w$. Let $S:=(v_{1},\ldots,v_{s})$ be the leaves of $T$ in the order they are encountered while traversing $C$. Subdivide each edge of $G$ by placing a vertex in the center of the edge and call the resulting graph $G^{\prime}$. Then $C$ is a proper-good curve for $G^{\prime}$ that contains $S$. Therefore $S$ is a free set in $G^{\prime}$. Therefore, for any $x_{1}<\cdots<x_{s}$ and any $y_{1},\ldots,y_{s}$, $G^{\prime}$ has a straight-line crossing-free drawing in which $v_{i}$ is placed at $(x_{i},y_{i})$ for each $i\in\{1,\ldots,s\}$. Any such drawing of $G^{\prime}$ gives a $1$-bend crossing-free drawing of $G$ in which $v_{i}$ is placed at $(x_{i},y_{i})$ for each $i\in\{1,\ldots,s\}$. Thus, $S$ is a $1$-bend collinear set in $G$. ∎
@@
The obvious question is whether the bound in 8 is of the correct form. Perhaps planar graphs have $1$-bend collinear sets of size $n-o(n)$. We can rule this possibility out with the following construction.
@@
The Goldner–Harary graph, $G_{0}$ is an $11$-vertex triangulation that is not Hamiltonian. We claim that any $1$-bend free set in $G_{0}$ has size at most $10$. Suppose, for the sake of contradiction, that $S:=V(G_{0})$ is a $1$-bend free set in $G_{0}$. Then $G$ has a $1$-bend crossing-free drawing with all vertices of $S$ placed on the ${x}$-axis. Since this is a $1$-bend crossing-free drawing and all vertices of $G_{0}$ are on the ${x}$-axis, no edge properly crosses the ${x}$-axis. Therefore, this is a $2$-page book embedding of $G_{0}$. This is impossible because any triangulation with a $2$-page book embedding is Hamiltonian.
For any positive integer $k$, let $G$ to be a graph consisting of $k$ disjoint copies of $G_{0}$. Then $G$ has $n=11k$ vertices. Any $1$-bend free set in $G$ must exclude one vertex from each copy of $G_{0}$ and therefore has size at most $1110k=10n/$ 11. ∎
@@ 6 Open Problems
    What is the largest value $\alpha$ such that every $n$-vertex planar graph has a free set of size $\Omega(n^{\alpha})$? Currently, we know that $1/2\leq\alpha\leq 0.9859$. Is $\alpha$ equal to the shortness exponent $\sigma$ for cubic triconnected planar graphs? In other words, does every triangulation $G$ contain a free set of size $\Omega(c(G^{*}))$?
@@
    Does every $n$-vertex planar graph of maximum degree $3$, $4$, $5$, or $6$ contain a free set of size $\Omega(n)$
@@
    Say that a subset $S$ of vertices in a planar graph $G$ is mostly free if, for every set $P$ of $|S|$ points in general position, there exists a straight-line crossing-free drawing of $G$ in which each vertex in $S$ is drawn at a distinct point in $P$. Is there a constant $\gamma>0$ such that every mostly free set $S$ contains a free subset of size at least $\gamma|S|$? The class of outerplanar graphs shows that $\gamma\leq 2/3$ since any set of $n$ points in general position is universal for the class of $n$-vertex outerplanar graphs [39, 15, 22] so the entire vertex set of every outerplanar graph is mostly free, but there exists $n$-vertex outerplanar graphs with no free set of size greater than $\lceil 2n\rceil$.
@@
  1. 5.
    Do universal point subset of size $\Omega(n^{1/2+\epsilon})$ for some $\epsilon>0$ exist for the class of $n$-vertex planar graphs. Currently, there is nothing that rules out a bound of $\Omega(n)$.
  2. 6.

### v2

@@ 5 A One-Bend Variant
A $k$-bend crossing-free drawing of a planar graph $G$ is a crossing-free drawing of $G$ in which each edge is represented by a polygonal chain consisting of at most $k+1$ line segments. Thus a 00-bend crossing-free drawing is a straight-line crossing-free drawing. This leads naturally to a definition of a $k$-bend free set introduced by Bose et al. [16]. An ordered set $S:=(v_{1},\ldots,v_{s})$ of vertices in a planar graph $G$ is a $k$-bend free set if, for any $x_{1}<\cdots<x_{s}$ and any $y_{1},\ldots,y_{s}$, there exists a $k$-bend crossing-free drawing $\Gamma$ of $G$ such that $(x_{i},y_{i})$ is the location of $v_{i}$ in $\Gamma$, for each $i\in\{1,\ldots,s\}$. As an application of 9, below, Bose et al. [16] prove the following result:
@@
They prove 8 by proving the following result, which is the main result in [16]:
@@
Refer to Fig. 12. Fix some straight-line crossing-free drawing of $G$, which also fixes a straight-line crossing-free drawing of $T$. Fatten the drawing of $T$ by some arbitrarily small value $\epsilon>0$ by taking the Minkowsky sum of $T$ with a disc of radius $\epsilon$. The boundary of this fattened tree is a simple closed curve $C_{0}$. Construct a curve $C$ by deforming $C_{0}$ in an $\epsilon$-neighbourhood of each leaf $v$ of $T$ so that it contains $v$. If $\epsilon$ is sufficiently small, then $C$ is a simple closed curve that intersects each edge $vw$ of $G$ in at most $2$ points: one point within distance $\epsilon$ of $v$ and one point within distance $\epsilon$ of $w$. Let $S:=(v_{1},\ldots,v_{s})$ be the leaves of $T$ in the order they are encountered while traversing $C$. Subdivide each edge of $G$ by placing a vertex in the center of the edge, and call the resulting graph $G^{\prime}$. Then $C$ is a proper-good curve for $G^{\prime}$ that contains $S$. Therefore, $S$ is a free set in $G^{\prime}$. Therefore, for any $x_{1}<\cdots<x_{s}$ and any $y_{1},\ldots,y_{s}$, $G^{\prime}$ has a straight-line crossing-free drawing in which $v_{i}$ is placed at $(x_{i},y_{i})$ for each $i\in\{1,\ldots,s\}$. Any such drawing of $G^{\prime}$ gives a $1$-bend crossing-free drawing of $G$ in which $v_{i}$ is placed at $(x_{i},y_{i})$ for each $i\in\{1,\ldots,s\}$. Thus $S$ is a $1$-bend collinear set in $G$. ∎
@@
The obvious question is whether the bound in 8 is of the correct form. Perhaps planar graphs have $1$-bend collinear sets of size $n-o(n)$. We can rule out this possibility with the following construction.
@@
The Goldner–Harary graph [40], $G_{0}$ is an $11$-vertex triangulation that is not Hamiltonian. We claim that any $1$-bend free set in $G_{0}$ has size at most $10$. Suppose, for the sake of contradiction, that $S:=V(G_{0})$ is a $1$-bend free set in $G_{0}$. Then $G_{0}$ has a $1$-bend crossing-free drawing $D$ with all vertices of $S$ placed on the ${x}$-axis. Since $D$ is a $1$-bend crossing-free drawing and all vertices of $G_{0}$ are on the ${x}$-axis, no edge properly crosses the ${x}$-axis.
We now show that the existence of such a crossing-free drawing $D$ implies that $G_{0}$ is Hamiltonian, contradicting our initial assumption. Let $v$ and $w$ be any two consecutive vertices of $G_{0}$ along the ${x}$-axis. Suppose, for contradiction, that $vw\not\in E(G_{0})$. Since a line segment between $v$ and $w$ does not cross any edge of $D$, adding the edge $vw$ to $G_{0}$ yields a planar graph with more edges than $G_{0}$. But this contradicts the fact that $G_{0}$ is a triangulation and therefore edge-maximal. Hence, every pair of consecutive vertices along the ${x}$-axis must be connected by an edge in $G_{0}$, implying that $G_{0}$ contains a Hamiltonian path. Similarly, a curve can be added to $D$ connecting the first and the last vertex on the ${x}$-axis without crossing any existing edge. An argument equivalent to the one above shows that this edge must also exist in $G_{0}$. It follows that $G_{0}$ contains a Hamiltonian cycle, completing the contradiction.
For any positive integer $k$, let $G$ be a graph consisting of $k$ disjoint copies of $G_{0}$. Then $G$ has $n=11k$ vertices. Any $1$-bend free set in $G$ must exclude one vertex from each copy of $G_{0}$ and therefore has size at most $1110k=10n/$ 11. ∎
@@ 6 Open Problems
    What is the largest value $\alpha$ such that every $n$-vertex planar graph has a free set of size $\Omega(n^{\alpha})$? Currently, we know that $1/2\leq\alpha\leq 0.9859$. Does $\alpha=\beta$, where $\beta$ is the largest value such that every $n$-vertex cubic triconnected planar graph contains a cycle of length $\Omega(n^{\beta})$?
@@
    Does every $n$-vertex planar graph of maximum degree $3$, $4$, $5$, or $6$ contain a free set of size $\Omega(n)$?
@@
    Say that a subset $S$ of vertices in a planar graph $G$ is mostly free if, for every set $P$ of $|S|$ points in general position, there exists a straight-line crossing-free drawing of $G$ in which each vertex in $S$ is drawn at a distinct point in $P$. Is there a constant $\gamma>0$ such that every mostly free set $S$ contains a free subset of size at least $\gamma|S|$? The class of outerplanar graphs shows that $\gamma\leq 2/3$ since any set of $n$ points in general position is universal for the class of $n$-vertex outerplanar graphs [41, 15, 22]. Thus the entire vertex set of every outerplanar graph is mostly free, but there exists $n$-vertex outerplanar graphs with no free set of size greater than $\lceil n/2\rceil$.
@@
  1. 7.
    Do universal point subset of size $\Omega(n^{1/2+\epsilon})$ for some $\epsilon>0$ exist for the class of $n$-vertex planar graphs. Currently, there is nothing that rules out a bound of $\Omega(n)$?
  2. 8.
@@
Acknowledgements
We thank Günter Rote and the anonymous referees for detailed and insightful comments, which significantly improved the quality of this manuscript.